Subject: APO 3.5m User's Committee meeting, 04/14/03


Submitted: Tue, 15 Apr 2003 12:01:04 -0400 (EDT)

Message number: 665 (previous: 664, next: 666 up: Index)

	  APO User's Committee Meeting, 04/14/03. 

Attending: Russet McMillan, John Holtzmann, Al Harper, Mike Shull, Bruce
Gillespie, Michael Strauss, Alan Uomoto, John Morse, Ed Turner, Bruce

Not able to attend: Chris Stubbs, Don York, Rene Walterbos

  Instrumentation news: DIS, Echelle, NIC-FPS
  Telescope news: Baffles, work on shutters
  Miscellaneous: Weather statistics, Board of Governor's meeting, etc. 

*******************Instrumentation news*****************

DIS installation of new red optics: The work on this is to start today
(4/14).  There is a danger that it will take a bit longer than the
alloted time, which means that people scheduled to use DIS in the days
following April 21 have a small danger of losing their time.

DIS blue scattering problem: As was reported at the last meeting, it
is now clear that there is a component of scattered light in the blue
camera, which now seems to have started when the blue side optics were
upgraded.  There is a guess that this is due to condensation on the
blue side.  We will pump on that side while the red installation is
going on, to see if this cleans it up.

New Echelle guider camera: This has just arrived on the mountain.
There are some things to be fixed (e.g., some of the cables are not
quite long enough), but this should be straightforward.  We hope to
install it next week, and put it on the sky around April 28.  At that
time, on-sky commissioning will take place.  It may take about 1/2
night; the goal is to get the system in place by the first echelle run
on May 10.  With this upgrade, the system should allow guiding on 17th
(and maybe even 18th) mag objects, opening up the echelle to quasar
work.  Note that most guiding is done off the slit; the field of view
of the guider is small enough (50 arcsec) that there is rarely a field
star bright enough to guide on.

Echelle ion pumps: These are showing an intermittent indication of
trouble.  They are working harder than usual and it is not clear why.
Over the last month, however, things have been looking better.  We have
suspicions of either outgassing, or perhaps simply aging of the
pumps (which in fact are nearing the end of their nominal lifetimes).
It may make sense to replace them during the summer shutdown; we hope
that there is no catastrophic failure before then.  

  There are also some concern of an increase in dark current as well,
which is due to some cross-talk between the chip and the ion pump; as
the latter works harder, the cross-talk increases.  Echelle users
perhaps should plan to take dark exposures before and after science
exposures to be careful.

  There was a formal review of NICFPS on April 4: Thanks to Bruce
Gillespie for putting together an excellent panel of seven people.
They heard presentations of the current status and future plans.
There will be a written report of recommendations.  The panel was
quite positive; it was clear that things were going well.  Jon
Holtzman had sent out a <a href=>request</a>
for input on the z-band filters (give URL); send Jon any comments you
might have.

********************Telescope and Enclosure news******************

  The tertiary baffle installation is now complete.  One concern was
that it would worsen the seeing (due to enclosing air in the tube of
the baffle), but we've seen some superb seeing nights since it has
been in.  But it won't be until the secondary baffle is also in place
(still several weeks away) that the scattered light situation will be
properly measured, but it should be better already. 

  On a related note: there are long-known asymmetries around the edge
of the primary that throw a few percent of the light of the PSF into
an extended halo, which can be seen around bright stars with SPICAM.
This was taken care of previously by simply masking (i.e., with
masking tape) the regions of the primary at fault.  However, this mask
was removed last August when the primary was realuminized, and so this
has been a problem ever since.  A more permanent solution, in terms of
an easily installed "cateye mask" is being built at Sunspot, and will
be ready in a few weeks.

  There is a 5-night bright time shutdown planned in May to replace
worn rollers in the shutter mechanism.  The telescope will be
completely unusable during that week.  This needs to be done in May,
not July; it requires having the dome open for an extended period, and
we want to avoid rain during this period.


Russet McMillan discussed weather statistics over the period September
1999 to March 2003; see the summary
at <a href=></a>.
  Overall, this past February was the worst ever (on the plus side,
all the snow received made this the first year of normal precipitation
in many years). 

  From Russet's statistics, a global average over several years has
shown that the telescope has been closed due to weather only 35% of
the time.  This of course is not the same as saying that it was
scientifically productive for every moment of the remaining 65%, but
nevertheless, this is a reassuring number. 

 The ARC Board of Governors will meet at APO on June 9 and 10, to
discuss the future of the SDSS and long-term plans for the 3.5m, and
the future of the ARC consortium in general.  Rene Walterbos will soon
send out the agenda for meeting, and points for discussion.  Each
institution can send up to 3 people to the meeting.  Each institution
should make sure to have a proper discussion internally of their
opinions on these issues.

  There have been several recent incidents whereby observers were
trying to arrange time swaps by going directly to the observing
specialists.  A reminder: such things should be done through your
institutional scheduler and the director.  There was confusion on
recent programs (especially transient programs): Who is observing?
Which instruments?  This has led to some wasted time.  One possibility
is to reinstate a program which was tried last fall: two half nights
were scheduled per week for transient science (there's always
*something* interesting going on.  This program was joint over a
number of institutions.

  Last month's minutes are approved. 

  The next meeting will be held on Monday, May 19, at 11:30 AM Eastern

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 665 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to